OpenStudy (anonymous):

8^(2x) =11 help please

OpenStudy (australopithecus):

Use logarithms Logrithmic rules 1) log_b(mn) = log_b(m) + logb(n) 2) log_b(m/n) = log_b(m) – logb(n) 3) log_b(m^(n)) = n · log_b(m) 4) log_3(3) = 1 5) log_e(x) = ln(x) 6) log(x) = y is the same as x =10^(y) 7) log_10(x) = log(x) 8) log(x^(2)) = 2log(x)

OpenStudy (australopithecus):

take the log of both sides log_8(8^(2x)) = log_8(11) Now use rules I just gave you

OpenStudy (australopithecus):

hint* use rule 8

OpenStudy (australopithecus):

then rule 4